# Measuring the Discrepancy between 3D Geometric Models using Directional Distance Fields

### v1

@@ I Introduction
Figure 1: Visual illustration of different distance metrics for 3D geometry data. For convenience, we use 2D illustration. The yellow points in (a), (b), and (c) refer to 3D points located on 3D surfaces indicated by curves, and the brown points in (d) represent the generated reference points. The blue arrows represent the established correspondence. The color in (d) changes from green to red indicating the distance fields of the two surfaces indicated by curves (i.e., the set of the distances of arbitrary points in 3D space to the surfaces).
@@
Existing distance metrics for 3D geometry data predominantly fall into two main categories, point-to-point (P2P) and point-to-face (P2F) distances. In the P2P category, two representative and commonly employed metrics are Earth Mover’s Distance (EMD) [14] and Chamfer Distance (CD) [15], as illustrated in Fig. 0(a) and 0(b), respectively. These metrics generally establish point-to-point correspondences between two point clouds [^1] Directly measuring the discrepancy between two continuous 3D surfaces presents challenges. When dealing with continuous 3D surfaces, such as triangle meshes, a typical sampling method is often applied to sample a set of points on the surface for computation. and subsequently compute the aggregation of the point-wise distances between corresponding points. However, these metrics suffer from limitations as they operate on sampled points, disregarding the continuity of the surfaces, which renders them ineffective. Additionally, the correspondence establishment process can be time-consuming, such as the bijection adopted in EMD. Alternatively, the P2F method [16], as illustrated in Fig. 0(c), mitigates the shortcomings of P2F to some extent, which involves sampling points from one surface and then computing each sampled point’s closest distance to the other surface to measure the discrepancy between two surfaces. However, the closest point searching could make it prone to falling into suboptimal solutions. Although several improved methods [17, 18, 19, 20, 21, 3] have been proposed, they still suffer from inefficiency or ineffectiveness. See the detailed review in Sec. II.
@@
To address this fundamental and challenging issue of quantifying the discrepancy between 3D geometric models, we propose a novel robust, efficient, and effective distance metric called DirDist.Distinguished from existing metrics, DirDist emphasizes the implicit representation of 3D geometric models, as depicted in Fig. 0(d). Methodologically, for any pair of 3D models, we begin by constructing their individual implicit fields, referred to as directional distance fields (DDFs). The discrepancy between these fields serves as an indicator of the distance between the corresponding 3D geometric models. To compute DirDist, we generate a set of reference points that are distributed near the models and shared by both models. For each reference point, we calculate its directional distances to the two models, capturing the local surface geometry. By taking the weighted average of the directional distance discrepancies contributed by all reference points, we obtain the final DirDist. The use of DDFs enables DirDist to robustly handle optimization convergence issues by effectively capturing surface geometry. Moreover, DirDist does not require a direct correspondence establishment process, resulting in high efficiency.
@@ III-B Problem Statement and Motivation
In what follows, we will detail the generation of reference points in Sec. III-C and the definition of DDF in Sec. III-D, leading to our efficient yet effective distance metric for 3D geometry data named DirDist in Sec. III-E. Finally, we also prove the essential properties of DirDist in Sec. III-F.
@@ III-C Generation of Reference Points
In the optimization/learning process of a typical geometric modeling task, it is common for one 3D model to remain unchanged as the ground truth, while the other one is optimized/learned. Without loss of generality, let’s assume $\mathcal{S}_{1}$ is the unchanged 3D model. In this scenario, we pre-define $\mathbf{Q}$ based on $\mathcal{S}_{1}$. Moreover, as $\mathbf{Q}$ are employed to induce the implicit distance field for the 3D surface, they should be distributed near the surface for meaningful distances.
@@
Specifically, we add Gaussian noise with a standard deviation $\sigma$ to the points on $\mathcal{S}_{1}$, displacing them away from the surface, as shown in Fig. 3. If $\mathcal{S}_{1}$ is represented in the form of a 3D point cloud, we directly introduce the Gaussian noise to its points and iterate this noise addition process multi times randomly, resulting in $M$ reference points situated closely to the surface. If $\mathcal{S}_{1}$ is represented with a triangle mesh, we first sample points on it to convert it to a point cloud, and then we generate $M$ reference points from the sampled point cloud through the aforementioned operation.
@@ III-D Directional Distance Field
Point Cloud. Let $\mathbf{P}:=\{\mathbf{p}\in\mathbb{R}^{3}\}_{i=1}^{N}$ be the point cloud representing the 3D model of $\mathcal{S}$ and $\Omega(\mathbf{q},\mathbf{P}):=\{\mathbf{p}_{k}\}_{k=1}^{K}$ the set of $K$-NN ($K$ Nearest Neighbor) points of $\mathbf{P}$ to $\mathbf{q}$. As shown in Fig. 3(a), $\mathbf{\hat{q}}$ must lie in the area that $\Omega(\mathbf{q},\mathbf{P})$ covers, and we thus approximate $\mathbf{\hat{q}}$ with the weighted averaging of the points in $\Omega(\mathbf{q},\mathbf{P})$:
@@ III-E DDF-based Distance Metric
III-F Proof of Essential Properties
The proposed DirDist in Eq. (7) possesses essential properties that align with standard distance metrics, including non-negativity, symmetry, identity of indiscernibles, and triangle inequality. Obviously, the first three properties are satisfied once the reference points are given. In the following, we show that it also fulfills the triangle inequality.
Specifically, we state and prove the following theorem:
Theorem 1. Given three surfaces, denoted as $\mathcal{S}_{1}$, $\mathcal{S}_{2}$, and $\mathcal{S}_{3}$, along with the generated reference point set $\mathbf{Q}$, the following inequality holds
@@ III-F Proof of Essential Properties
For each $\mathbf{q}\in\mathbf{Q}$, we denote $d(\mathbf{q},\mathcal{S}_{1},\mathcal{S}_{3})$, $d(\mathbf{q},\mathcal{S}_{1},\mathcal{S}_{2})$, and $d(\mathbf{q},\mathcal{S}_{2},\mathcal{S}_{3})$ as $d_{13}$, $d_{12}$, and $d_{23}$, respectively. According to the triangle inequality, they satisfies
@@
As describe in Sec. 3, the reference points $\mathbf{Q}$ are generated at the area near the surfaces, thus, the differences introduced by the reference points are sufficiently small, such that they are less than $\frac{1}{\beta}$, i.e., $d_{13},d_{12},d_{23}<\frac{1}{\beta}$.
@@ IV-A Template Surface Fitting
Here, we consider the recent pipeline introduced in [52, 53]. Specifically, in order to avoid face intersections during the surface deformation, we optimize the diffusion reparameterization rather than the vertices of $\mathcal{S}_{\rm init}$. In this process, the distance between the deformed initial surface and the target surface is used to guide the optimization, which is critical in the whole progress. Technically, the diffusion reparameterization is defined as $\mathbf{u}=(\mathbf{I}+\alpha\mathbf{L})\mathbf{V}$, where $\mathbf{L}\in\mathbb{R}^{N\times N}$ is the discrete Laplace operator, $\mathbf{I}\in\mathbb{R}^{N\times N}$ refers to the identity matrix, and $\alpha$ is a constant weight. Consequently, the deformed surface, denoted as $\mathcal{S}^{{}^{\prime}}$, possesses vertices $\mathbf{V}^{\prime}=(\mathbf{I}+\alpha\mathbf{L})^{-1}\mathbf{u}$, while retaining the same face indices as $\mathcal{S}_{\rm init}$. $\mathbf{u}$ could be optimized by minimizing the distance between $\mathcal{S}^{{}^{\prime}}$ and $\mathcal{S}_{\rm tgt}$:
@@ IV-B Rigid Registration of 3D Point Clouds
Data: The source surface with vertex set $\mathbf{V}$ and face index set $\mathbf{F}$; the distance threshold $\epsilon$.
Result: The coordinates of deformation nodes $\mathbf{V}_{\rm DF}$.
@@ IV-E Human Pose Optimization from Point Clouds
In the domain of digital human research, the 3D data representing humans are usually collected through some common 3D scanners, e.g. RGBD cameras and Lidars. As shown in Fig. 7, given the initial human pose, which differs from the scanned data, we need to refine it according to the scanned data, making it aligned with the scanned data. Skinned Multi-Person Linear Model (SMPL) [55] is the most widely used parametric model to represent the 3D human body. Specifically, the SMPL model is a triangle mesh with the human body shape, and its pose is controlled by a matrix $\mathbf{J}\in\mathbb{R}^{24\times 3}$, thus, various values of $\mathbf{J}$ could represent the human models with different pose. In reality, we need to optimize the human pose according to the given human scan, which is usually a whole or partial point cloud. This could be achieved by minimizing the difference between the SMPL model and the scanned point cloud. Denote by $\mathbf{P}_{\rm scan}$ the scanned point cloud, the human pose estimation could be formulated as
@@ V-A2 Comparisons
We chose two baseline distance metrics for comparison, i.e., CD and P2F. For a fair comparison, the numbers of sampled points for computing CD and P2F were kept the same as our DirDist. Besides, we also made a comparison with MDA [53], where a one-sided CD is used as the distance metric and the other settings were kept the same as their original paper. We utilized Normal Consistency (NC) and F-Score with thresholds of 0.5% and 1%, denoted as F-0.005 and F-0.01, as the evaluation metrics. We refer the readers to [28] for the detailed definitions. During the evaluation, 500K points were sampled on the deformed and target surfaces. Table I and Fig. 8 show the numerical and visual results, respectively, where both quantitative accuracy and visual quality of deformed shapes with our DirDist are much better than baseline methods. While MDA exhibits superior visual results compared to CD and P2F, its numerical results about F-0.005 and F-0.01 in Table I is worse than both. This discrepancy arises from the fact that the deformed surfaces produced by MDA are excessively smoothed, leading to a convergence to local optima, because of its loss function. Although the overall shape is similar, this oversmoothing effect is evident in the error maps illustrated in Fig. 9.
@@
V-B Rigid Registration of 3D Point Clouds
V-B1 Implementation Details
We experimented with the commonly used 3DMatch dataset [59]. Following previous works [60, 61], we downsampled the original point clouds through Voxel Grid Filtering with a resolution of 5cm, producing point clouds each with 5K points. Then we chose a feature-based registration methods, MAC [62], to calculate the coarse transformation between the two point clouds, serving as the initialization of the optimization of Eq. (13).To test the robustness of different distance metrics, we also conducted experiments on the point clouds with noise and outliers. To simulate the noisy conditions, we added Gaussian noise with a mean of zero and a standard deviation of 2cm to the clean point clouds. For the outlier data, we infused the point clouds with an additional 50% randomly selected points. In all registration experiments, we set $M$ 10 times the number of points, $K=5$, $\beta=20$, and $\sigma=0.05$. We employed the Adam optimizer to optimize, spanning 200 iterations with a learning rate set at 0.02.
TABLE II: Quantitative comparisons of rigid registration on the 3DMatch dataset [59], where the coarse transformation is estimated through MAC [62].

### v4

# DirDist: A Metric for Comparing 3D Shapes Using Directional Distance Fields

@@ I Introduction
Figure 1: Visual illustration of different distance metrics for 3D geometry data. For convenience, we use 2D illustration. The yellow points in (a), (b), (c), and (d) refer to 3D points located on 3D surfaces indicated by curves, and the brown points in (e) represent the generated reference points. The blue arrows represent the established correspondence. The color in (e) changes from green to red indicating the distance fields of the two surfaces indicated by curves (i.e., the set of the distances of arbitrary points in 3D space to the surfaces).
@@
Existing distance metrics for 3D geometry data predominantly fall into two main categories, point-to-point (P2P) and point-to-face (P2F) distances. In the P2P category, two representative and commonly employed metrics are Earth Mover’s Distance (EMD) [14] and Chamfer Distance (CD) [15], as illustrated in Figs. 1(a) and 1(b), respectively. These metrics generally establish point-to-point correspondences between two point clouds [^1] Directly measuring the discrepancy between two continuous 3D surfaces presents challenges. When dealing with continuous 3D surfaces, such as triangle meshes, a typical sampling method is often applied to sample a set of points on the surface for computation. and subsequently compute the aggregation of the point-wise distances between the corresponding points. However, these metrics suffer from limitations as they operate on sampled points, disregarding the continuity of the surfaces, which renders them ineffective. Additionally, the correspondence establishment process can be time-consuming, such as the bijection adopted in EMD. Alternatively, the P2F method [16], as illustrated in Fig. 1(c), mitigates the shortcomings of P2F to some extent, which involves sampling points from one surface and then computing each sampled point’s closest distance to the other surface to measure the discrepancy between two surfaces. However, the closest point search could make it prone to suboptimal solutions. In contrast, the ARL method [3] employs randomly sampled lines to intersect with the two surfaces, utilizing the resulting intersection points to establish correspondences, as illustrated in Fig. 1(d). However, the positions of these intersection points are inherently dependent on the spatial arrangement of the sampled lines. Consequently, the use of randomly sampled lines may introduce inaccuracies in the correspondence between the geometric data, potentially leading to erroneous alignments. Although several improved methods [17, 18, 19, 20, 21] have been proposed, they still suffer from inefficiency or ineffectiveness. See the detailed review in Sec. II.
@@
To address this fundamental and challenging issue of quantifying the discrepancy between 3D geometric models, we propose a novel robust, efficient, and effective distance metric called DirDist.Distinguished from existing metrics, DirDist emphasizes the implicit representation of 3D geometric models, as depicted in Fig. 1(e). Methodologically, for any pair of 3D models, we begin by constructing their individual implicit fields, referred to as directional distance fields (DDFs). The discrepancy between these fields serves as an indicator of the distance between the corresponding 3D geometric models. To compute DirDist, we generate a set of reference points that are distributed near the models and shared by both models. For each reference point, we calculate its directional distances to the two models, capturing the local surface geometry. By taking the weighted average of the directional distance discrepancies contributed by all reference points, we obtain the final DirDist. The use of DDFs enables DirDist to robustly handle optimization convergence issues by effectively capturing surface geometry. Moreover, DirDist does not require a direct correspondence establishment process, resulting in high efficiency.
@@ III-B Problem Statement and Motivation
In what follows, we will detail the generation of reference points in Sec. III-C and the definition of DDF in Sec. III-D, leading to our efficient yet effective distance metric for 3D geometry data named DirDist in Sec. III-E.
@@ III-C Generation of Reference Points
In the optimization/learning process of a typical geometric modeling task, it is common for one 3D model to remain unchanged as the ground truth, while the other one is optimized/learned such that it closely approximates the ground truth. Consequently, we focuses on the regions near the unchanged model, and the reference points should be generated within these regions.
@@
Without loss of generality, let’s assume $\mathcal{S}_{1}$ is the unchanged 3D model. In this scenario, we pre-define $\mathbf{Q}$ based on $\mathcal{S}_{1}$. Specifically, we add Gaussian noise with a standard deviation $\sigma$ to the points on $\mathcal{S}_{1}$, displacing them away from the surface, as shown in Fig. 3. If $\mathcal{S}_{1}$ is represented in the form of a 3D point cloud, we directly introduce the Gaussian noise to its points and iterate this noise addition process multi times randomly, resulting in $M$ reference points situated closely to the surface. If $\mathcal{S}_{1}$ is represented with a triangle mesh, we first sample points on it to convert it to a point cloud, and then we generate $M$ reference points from the sampled point cloud through the aforementioned operation.
@@ III-D Directional Distance Field
Point Cloud. Let $\mathbf{P}:=\{\mathbf{p}_{i}\in\mathbb{R}^{3}\}_{i=1}^{N}$ be the point cloud representing the 3D model of $\mathcal{S}$ and $\Omega(\mathbf{q},\mathbf{P}):=\{\mathbf{p}_{k}\}_{k=1}^{K}$ the set of $K$-NN ($K$ Nearest Neighbor) points of $\mathbf{P}$ to $\mathbf{q}$. As shown in Fig. 4(a), $\mathbf{\hat{q}}$ must lie in the area that $\Omega(\mathbf{q},\mathbf{P})$ covers, and we thus approximate $\mathbf{\hat{q}}$ with the weighted averaging of the points in $\Omega(\mathbf{q},\mathbf{P})$:
@@ III-E DDF-based Distance Metric
The proposed DirDist in Eq. (7) possesses essential properties that align with standard distance metrics, including non-negativity, symmetry, the identity of indiscernibles, and the triangle inequality. Obviously, the first three properties are satisfied once the reference points are given. Theorem 1 shows that it also satisfies the triangle inequality. In addition, under some specific settings, existing distance metrics, such as CD and P2F, can be regarded as special cases of our DirDist. Theorems 2 and 3 illustrate these properties.
Theorem 1.
Given three surfaces, denoted as $\mathcal{S}_{1}$, $\mathcal{S}_{2}$, and $\mathcal{S}_{3}$, along with the generated reference point set $\mathbf{Q}$, the following inequality holds
@@
For each $\mathbf{q}\in\mathbf{Q}$, we denote $d(\mathbf{q},\mathcal{S}_{1},\mathcal{S}_{3})$, $d(\mathbf{q},\mathcal{S}_{1},\mathcal{S}_{2})$, and $d(\mathbf{q},\mathcal{S}_{2},\mathcal{S}_{3})$ as $d_{13}$, $d_{12}$, and $d_{23}$, respectively. According to the triangle inequality, they satisfy
@@
As described in Sec. III-C of the manuscript, the reference points $\mathbf{Q}$ are generated at the area near the surfaces, thus, the differences introduced by the reference points are sufficiently small, such that they are less than $\frac{1}{\beta}$, i.e., $d_{13},d_{12},d_{23}<\frac{1}{\beta}$.
@@
Theorem 2.
Given two point clouds, $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$, when the reference points are defined as $\mathbf{Q}=\mathbf{P}_{1}\cup\mathbf{P}_{2}$, and the DDF is defined as $\mathcal{F}=f$ with parameters $K=1$ and $\beta=0$, DirDist becomes equivalent to CD.
Proof.
Let first consider the reference points from $\mathbf{q}\in\mathbf{P}_{1}$, according to the definition of DDF, their DDFs of two point clouds can be calculated through
$\mathcal{F}_{\mathbf{P}_{1}}(\mathbf{q})=f_{\mathbf{P}_{1}}(\mathbf{q})=\|\texttt{NN}({\mathbf{q}},\mathbf{P}_{1})-\mathbf{q}\|_{2}=\|\mathbf{q}-\mathbf{q}\|=0,$
$\mathcal{F}_{\mathbf{P}_{2}}(\mathbf{q})=f_{\mathbf{P}_{2}}(\mathbf{q})=\|\texttt{NN}(\mathbf{q},\mathbf{P}_{2})-\mathbf{q}\|_{2},$
where NN is the 1-NN operation. Thus, DirDist under these reference points is
$\mathcal{D}_{1}$ $=\sum_{\mathbf{q}\in\mathbf{P}_{1}}|\mathcal{F}_{\mathbf{P}_{1}}(\mathbf{q})-\mathcal{F}_{\mathbf{P}_{2}}(\mathbf{q})|$
$=\sum_{\mathbf{q}\in\mathbf{P}_{1}}|0-\|\texttt{NN}(\mathbf{q},\mathbf{P}_{2})-\mathbf{q}\|_{2}|$
$=\sum_{\mathbf{q}\in\mathbf{P}_{1}}\|\texttt{NN}(\mathbf{q},\mathbf{P}_{2})-\mathbf{q}\|_{2}.$
Similarity, DirDist under the reference points $\mathbf{q}\in\mathbf{P}_{2}$ is
$\mathcal{D}_{2}=\sum_{\mathbf{q}\in\mathbf{P}_{2}}\|\texttt{NN}(\mathbf{q},\mathbf{P}_{1})-\mathbf{q}\|_{2}.$
Obviously, $\mathcal{D}_{\rm DirDist}=\mathcal{D}_{1}+\mathcal{D}_{2}$ is the same as CD.
Theorem 3.
Given two triangle meshes, $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, when the reference points $\mathbf{Q}=\{\mathbf{q}_{i}\}_{i=1}^{M}$ are sampled on the surface, and the DDF is defined as $\mathcal{F}=f$ with parameter $\beta=0$, DirDist becomes equivalent to the P2F distance.
Proof.
Let first considering the reference point sampled from $\mathcal{S}_{1}$, $\mathbf{Q}_{1}=\{\mathbf{q}_{i}\}_{i=1}^{M_{1}}$, according to the definition of DDF, the DDFs of two triangle meshes can be represented as
$\mathcal{F}_{\mathbf{P}_{1}}(\mathbf{q}_{i})=f_{\mathbf{P}_{1}}(\mathbf{q}_{i})=\|\texttt{NP}(\mathbf{q}_{i},\mathcal{S}_{1})-\mathbf{q}_{i}\|_{2}=\|\mathbf{q}_{i}-\mathbf{q}_{i}\|=0,$
$\mathcal{F}_{\mathbf{P}_{2}}(\mathbf{q}_{i})=f_{\mathbf{P}_{2}}(\mathbf{q}_{i})=\|\texttt{NP}(\mathbf{q}_{i},\mathcal{S}_{2})-\mathbf{q}_{i}\|_{2},$
where NP is the nearest point of the given point on the surface. Thus, DirDist unde these reference points is
$\mathcal{D}_{1}$ $=\sum_{\mathbf{q}_{i}\in\mathbf{Q}_{1}}|\mathcal{F}_{\mathbf{P}_{1}}(\mathbf{q}_{i})-\mathcal{F}_{\mathbf{P}_{2}}(\mathbf{q}_{i})|$
$=\sum_{\mathbf{q}_{i}\in\mathbf{Q}_{1}}|0-\|\texttt{NP}(\mathbf{q}_{i},\mathcal{S}_{2})-\mathbf{q}_{i}\|_{2}|$
$=\sum_{\mathbf{q}_{i}\in\mathbf{Q}_{1}}\|\texttt{NP}(\mathbf{q}_{i},\mathcal{S}_{2})-\mathbf{q}_{i}\|_{2}$
Similarity, DirDist under the reference points sampled from $\mathcal{S}_{2}$, $\mathbf{Q}_{2}=\{\mathbf{q}_{i}\}_{i=1}^{M_{2}}$ is
$\mathcal{D}_{2}=\sum_{\mathbf{q}_{i}\in\mathbf{Q}_{2}}\|\texttt{NP}(\mathbf{q}_{i},\mathcal{S}_{1})-\mathbf{q}_{i}\|_{2}.$
Obviously, $\mathcal{D}_{\rm DirDist}=\mathcal{D}_{1}+\mathcal{D}_{2}$ is the same as P2F distance. ∎
@@ IV-A Template Surface Fitting
Here, we consider the recent pipeline introduced in [52, 53]. Specifically, in order to avoid face intersections during the surface deformation, we optimize the diffusion reparameterization rather than the vertices of $\mathcal{S}_{\rm init}$. In this process, the distance between the deformed initial surface and the target surface is used to guide the optimization, which is critical in the whole process. Technically, the diffusion reparameterization is defined as $\mathbf{u}=(\mathbf{I}+\alpha\mathbf{L})\mathbf{V}$, where $\mathbf{L}\in\mathbb{R}^{N\times N}$ is the discrete Laplace operator, $\mathbf{I}\in\mathbb{R}^{N\times N}$ refers to the identity matrix, and $\alpha$ is a constant weight. Consequently, the deformed surface, denoted as $\mathcal{S}^{{}^{\prime}}$, possesses vertices $\mathbf{V}^{\prime}=(\mathbf{I}+\alpha\mathbf{L})^{-1}\mathbf{u}$, while retaining the same face indices as $\mathcal{S}_{\rm init}$. $\mathbf{u}$ could be optimized by minimizing the distance between $\mathcal{S}^{{}^{\prime}}$ and $\mathcal{S}_{\rm tgt}$:
@@ IV-B Rigid Registration of 3D Point Clouds
1
Input: The source surface with vertex set $\mathbf{V}$ and face index set $\mathbf{F}$; the distance threshold $\epsilon$.
Output: The coordinates of deformation nodes $\mathbf{V}_{\rm DF}$.
@@ IV-E Human Pose Optimization from Point Clouds
In the domain of digital human research, the 3D data representing humans are usually collected through some common 3D scanners, e.g. RGBD cameras and Lidars. As shown in Fig. 7, given the initial human pose, which differs from the scanned data, we need to refine it according to the scanned data, making it aligned with the scanned data. Skinned Multi-Person Linear Model (SMPL) [55] is the most widely used parametric model to represent the 3D human body.
@@
Specifically, the SMPL model is a triangle mesh with the human body shape, and its pose is controlled by a matrix $\mathbf{J}\in\mathbb{R}^{24\times 3}$, thus, various values of $\mathbf{J}$ could represent the human models with different pose. In reality, we need to optimize the human pose according to the given human scan, which is usually a whole or partial point cloud. This could be achieved by minimizing the difference between the SMPL model and the scanned point cloud. Denote by $\mathbf{P}_{\rm scan}$ the scanned point cloud, the human pose estimation could be formulated as
@@ V-A2 Comparisons
We chose two baseline distance metrics for comparison, i.e., CD and P2F. For a fair comparison, the numbers of sampled points for computing CD and P2F were kept the same as our DirDist. Besides, we also made a comparison with MDA [53], where a one-sided CD is used as the distance metric and the other settings were kept the same as their original paper. We utilized Normal Consistency (NC) and F-Score with thresholds of 0.5% and 1%, denoted as F-0.005 and F-0.01, as the evaluation metrics. We refer the readers to [28] for the detailed definitions. During the evaluation, 500K points were sampled on the deformed and target surfaces. Table I and Fig. 8 show the numerical and visual results, respectively, where both quantitative accuracy and visual quality of deformed shapes with our DirDist are much better than baseline methods. While MDA exhibits superior visual results compared to CD and P2F, its numerical results about F-0.005 and F-0.01 in Table I is worse than both. This discrepancy arises from the fact that the deformed surfaces produced by MDA are excessively smoothed, leading to a convergence to local optima, because of its loss function. Although the overall shape is similar, this oversmoothing effect is evident in the error maps illustrated in Fig. 9. Table II presents the running time and GPU memory costs of different distance metrics per iteration. Notably, our DirDist incurs a computational cost similar to P2F but higher than CD. This is because both DirDist and P2F require identifying the closest points on triangle meshes, a more complex task compared to CD, which only involves finding the nearest points in point clouds.
@@
TABLE II: Running time and GPU memory costs of different distance metrics in the template surface fitting task.
Method CD P2F Ours
Running Time (ms/Iter.) 49 161 159
GPU Memory (MB) 873 891 893
TABLE III: Quantitative comparisons of rigid registration on the 3DMatch dataset [59], where the coarse transformation is estimated through MAC [60].
@@
[…]
V-B Rigid Registration of 3D Point Clouds
V-B1 Implementation Details
We experimented with the commonly used 3DMatch dataset [59]. Following previous works [61, 62], we downsampled the original point clouds through Voxel Grid Filtering with a resolution of 5cm, producing point clouds each with approximately 5K points. Then we chose a feature-based registration methods, MAC [60], to calculate the coarse transformation between the two point clouds, serving as the initialization of the optimization of Eq. (13). To test the robustness of different distance metrics, we also conducted experiments on the point clouds with noise and outliers. To simulate the noisy conditions, we added Gaussian noise with a mean of zero and a standard deviation of 2cm to the clean point clouds. For the outlier data, we infused the point clouds with an additional 50% randomly selected points. In all registration experiments, we set $M$ 10 times the number of points, $K=5$, $\beta=20$, and $\sigma=0.05$. We employed the Adam optimizer to optimize, spanning 200 iterations with a learning rate set at 0.02.